OpenStudy (anonymous):

What would you have to invest today at 5% interest compounded monthly if you wish to have $20,000 in 10 years?

Parth (parthkohli):

\[P\left(1 + \dfrac{5}{12\cdot 100}\right)^{10 \cdot 12} = 20000\]Solve for \(P\)

OpenStudy (anonymous):

What should i do if my calulater keeps saying ERR: Overflow

Parth (parthkohli):

OK, let me explain. You must divide both sides by \(\left(1 + \dfrac{5}{12 \cdot 100}\right)^{10 \cdot 12}\).

Parth (parthkohli):

\[P\left(1 + \dfrac{1}{240}\right)^{120} = 20000\]

OpenStudy (anonymous):

Order of operations.

Parth (parthkohli):

OK wait, is that 5% per month or per annum?

Parth (parthkohli):

Wait a second.

Parth (parthkohli):

\[P\left(1 + \dfrac{5}{100}\right)^{120} = 20000\]If that is 5% per month, then it is a little easier...

OpenStudy (anonymous):

and where did you get 100 from should it no be 10 because when i look back at the book she claims that you would do 12 for the number of months in a year and there are 10 years

Parth (parthkohli):

The formula for compound interest is,\[P \left(1 + \dfrac{r}{100}\right)^n\]

Parth (parthkohli):

\(n\) is the number of times it is compounded, which would be \(12 \times 10 = 120\) and \(r\) is the rate which is \(5\). But does the amount compound monthly or yearly?

OpenStudy (anonymous):

monthly and my book say r over n not 100

Parth (parthkohli):

Hmm, don't get confused with that. Our books mean different stuff. Your book means \(r\%\) by \(r\).

Parth (parthkohli):

Oh, I meant to ask if it was 5% annually or monthly.

Parth (parthkohli):

I'd assume that it means annually.\[P\left(1 + \dfrac{5}{100 \cdot 12}\right)^{12 \times 10} = 20,000\]

OpenStudy (anonymous):

its monthly and the is my formula A=P(1+ R/N) raise NT (sorry i cant do the upsidedown V)